OpenStudy (bananas1234):

if f(x) = 9 - 4/x + 1/, f(-5)

OpenStudy (zeesbrat3):

Just plug -5 in for x

OpenStudy (zeesbrat3):

\[f(-5) = 9 - \frac{ 4 }{ (-5) }+1\]

OpenStudy (zeesbrat3):

\[f(-5) = 9 + \frac{ 4 }{ 5 }+1\]

OpenStudy (zeesbrat3):

Do you understand?

OpenStudy (bananas1234):

i think so, what is the next step? @zeesbrat3

OpenStudy (zeesbrat3):

Add them all together

OpenStudy (bananas1234):

how do i do that when its a fraction? @zeesbrat3

OpenStudy (zeesbrat3):

Well there are two ways of doing it. 1. Find the decimal version of 4/5 2. Change the numbers so that they are in fraction form as well and add the numerators

OpenStudy (zeesbrat3):

The first way: \[f(-5) = 9 + .8 +1\] The second way: \[f(-5) = \frac{ 45 }{ 5 } + \frac{ 4 }{ 5 } + \frac{ 5 }{ 5 }\]

OpenStudy (bananas1234):

that would make the answer 10.8 and that is not an option.

OpenStudy (zeesbrat3):

What are you options?

OpenStudy (bananas1234):

-15 -7 7 23 I think we are looking at this the wrong way

OpenStudy (zeesbrat3):

Are you sure the expression is f(x) = 9 + 4/x +1?

OpenStudy (bananas1234):

no. its \[if f(x) = 9 - 4|x + 1| find f(-5)\]

OpenStudy (zeesbrat3):

That would make a lot more sense

OpenStudy (zeesbrat3):

\[f(x) = 9 - 4\left| x +1\right|\] \[f(-5) = 9 - 4\left| (-5) + 1) \right|\] \[f(-5) = 9 -4\left| -4 \right|\] \[f(-5) = 9 - 4(4)\] \[f(-5) = 9 -16\] \[f(-5) = -7\]

OpenStudy (bananas1234):

that makes sense

OpenStudy (zeesbrat3):

Does it, or are you just saying that? Feel free to ask questions if you have them.

OpenStudy (mathmale):

Better check that you have copied this problem down correctly. You've written: f(x) = 9 - 4/x + 1/ Do you mean \[f(x)=9-\frac{ 4 }{ x } or \]

OpenStudy (mathmale):

Better check that you have copied this problem down correctly. You've written: f(x) = 9 - 4/x + 1/ Do you mean \[f(x)=9-\frac{ 4 }{ x } + 1, ~or~\frac{ 9-4 }{ x+1 },~or~... ~?\]

OpenStudy (mathmale):

Please use parentheses to indicate which work has to be done first (following order of operations rules).